geerky42 (geerky42):

A fair six-sided dice with sides labeled 1, 1, 2, 3, 3, 6 is thrown. What is the expected value of a single roll?

OpenStudy (anonymous):

(1/6)(1) + (1/6)(1) + (1/6)(2) + (1/6)(3) + (1/6)(3) + (1/6)(6) = (1/6)(1 + 1 + 2 + 3 + 3 + 6) = (1/6)(16) = 8/3 That is determined basically by adding the probability of each value by the actual value. Notice how 6(1/6) = 1, so all possibilities are taken into account.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @geerky42